AN ADMINISTRATOR OF A POPULAR WEB SITE IS TOLD THAT A NEW SERVER CAN HANDLE 11,000 HITS (USERS ACCESSING THE SITE) PER SECOND. THE WEB SITE CURRENTLY EXPERIENCES A PEAK DEMAND OF ABOUT 85,000 HITS PER SECOND; BUT EVERY MONTH THE PEAK DEMAND INCREASES BY 3500 HITS PER SECOND. USE A PROPORTION EQUATION TO DETERMINE HOW MANY NEW SERVERS THE ADMINISTRATOR SHOULD BUY TO ADDRESS EXPECTED TRAFFIC FOR THE NEXT 18 MONTHS.
Answer by stanbon(75887)

AN ADMINISTRATOR OF A POPULAR WEB SITE IS TOLD THAT A NEW SERVER CAN HANDLE 11,000 HITS (USERS ACCESSING THE SITE) PER SECOND. THE WEB SITE CURRENTLY EXPERIENCES A PEAK DEMAND OF ABOUT 85,000 HITS PER SECOND; BUT EVERY MONTH THE PEAK DEMAND INCREASES BY 3500 HITS PER SECOND. USE A PROPORTION EQUATION TO DETERMINE HOW MANY NEW SERVERS THE ADMINISTRATOR SHOULD BUY TO ADDRESS EXPECTED TRAFFIC FOR THE NEXT 18 MONTHS.
------------------------------
If the current number of servers is handlng 85000 hit at a rate of 11000
hits per server, there must be more than 85000/11000 servers on line at
present; or > 8.
-----------------
Using 8 as the present # of servers set up the proportion:
x/[85000+18*3500] = 8/85000
x/148000 = 8/85000
x = 148000*(8/85000)
x = 13.93
---------------
They will need at least 14 servers in 18 months, or 6 additional ones.
=================
Cheers,
Stan H.